the length of a rectangle is 12 cm more than twice its width. find the largest possible width if the perimeter is at most 120cm
Answer by richwmiller(17219)

L=2W+12
2L+2W=120
2(2W+12)+2W=120
4W+24+2W=120
6W=96
W=16 max